Write the formula needed


Case 1.1 (Moving observer)

1) Frequency Heard increases ( Towards)

2) Frequency heard decreases (Away)


Case1.2 (Moving sourcer)

3) Frequency heard fades (Move away)

4) Frequency heard increases (Towards)


Case 2. The Source is moving relative to a stationary observer.

5) Moving towards each other equation

6) Moving away from each other equation

7) wavelength Formula

8) Sound speed in air (V)

9) Beat Frequency

10) Perceived beat frequency on observer ear

11) Density volume of regular object

12) Specific Gravity


1
Expert's answer
2021-02-18T18:55:58-0500

1)


"f_o=f_s\frac{v+v_o}{v}"

2)


"f_o=f_s\frac{v-v_o}{v}"

3)


"f_o=f_s\frac{v}{v+v_s}"

4)


"f_o=f_s\frac{v}{v-v_s}"

5)


"V=v+u"

6)


"V=v-u"

7)


"\lambda=\frac{v}{f}"

8)


"v=\sqrt{\frac{\gamma RT}{M}}"

9)


"f_b=f_2-f_1"

10)


"f_p=0.5(f_2+f_1)"

11)


"\rho=\frac{m}{V}"

12)


"SG=\frac{\rho}{\rho_w}"